# Optimal Methods for Unknown Piecewise Smooth Problems I: Convex Optimization

### v1

@@ 3 Accelerated Prox-level method for piecewise smooth optimization
Note that the optimal solution $x^{*}$ is always feasible for the subproblem if we set the level $\tilde{l}=f^{*}$. Hence, any feasible solution to the subproblem $x^{t}$ should satisfy $\left\|x^{t}-\bar{y}\right\|\leq\left\|x^{*}-\bar{y}\right\|$. ∎∎
One remaining question is that, even though $\bar{L}\left(t\right)$ (the average of $L_{t}$, defined in (8)) is computable in practice, it is still unclear how to bound the constant $\bar{L}\left(t\right)$ as it appears in Proposition 3.1 and Theorem 3.1. In the following Proposition 3.2, we show that value $\bar{L}\left(t\right)$ can be upper bounded by a weighted average of empirical Lipschitz smoothness constants $\tilde{L}_{t}$, where $\tilde{L}_{t}$ is the best Lipschitz constant between $\bar{x}^{t,r}$ and $\underline{x}^{t,l+1}$ with $r>l$ in $t$-th outer iteration.
@@
then $L_{t}\leq\tilde{L}_{t}$ ($L_{t}$ is defined in (8)). Furthermore, let the weighted sequence $\omega_{t}$ satisfy $\omega_{t}=\omega_{t-1}\left(1-{\alpha_{t}}/{2}\right)^{-1}$. Then $\bar{L}\left(N\right)$ (defined in (8)) admits the upper bound:
$\bar{L}\left(N\right)<\frac{3\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\Big/\left(\alpha_{t}\cdot\tilde{L}_{t}\right)}\ .$ (18)
@@
where $(a)$ holds by the definition of $L_{t}$ and $(b)$ follows from $f\left(\hat{x}^{t}\right)\leq f\left(\hat{x}^{t-1}\right)$ and the definition of $\mathcal{E}(t)$. Finally, since $L_{t}\leq\tilde{L}_{t}$, the stated upper bound follows. ∎∎
@@ 4 Normalized Wolfe certificate
Definition 4.1 (Normalized Wolfe certificate, W-certificate).
@@
$\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)=\frac{1}{\iota}\max_{x\in\mathcal{B}\left(\bar{y},\iota\right)\cap X}\left(\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right)\leq\nu\ ,\ \text{where}\ \psi_{\mathcal{P},f}\left(x\right)=\max_{z\in\mathcal{P}}\left\{\ell_{f}\left(x,z\right)\right\}\ .$ (19)
@@
$\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ (defined in (19)) is non-increasing with respect to $\iota$.
@@
$\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\left\|\tilde{x}-\bar{y}\right\|}\geq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\left\|x_{\iota 1}-\bar{y}\right\|}\ .$ (20)
Combining (20), $\left\|\bar{y}-\tilde{x}\right\|=\frac{\iota_{2}}{\iota_{1}}\left\|x_{\iota 1}-\bar{y}\right\|$ gives $\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\iota_{2}}\geq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\iota_{1}}$. By the definition of $x_{\iota 2}$, we have $\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\iota_{1}}\leq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\iota_{2}}\leq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 2}\right)}{\iota_{2}}$. Hence, we complete our proof of that $\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ is non-increasing with respect to $\iota$. ∎∎
@@
$\begin{split}f\left(\bar{y}\right)-f^{*}&\leq\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{p}^{*}\right)\\ &\leq\max_{x\in\mathcal{B}\left(\bar{y},\left\|x_{p}^{*}-\bar{y}\right\|\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &=\left\|x_{p}^{*}-\bar{y}\right\|\mathcal{V}_{\mathcal{P},f}\left(\left\|x_{p}^{*}-\bar{y}\right\|;\bar{y}\right)\\ &\leq\left\|x_{p}^{*}-\bar{y}\right\|\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)\leq\left\|x_{p}^{*}-\bar{y}\right\|\nu\end{split}\ .$ (21)
@@
$\left\|x_{p}^{*}-\bar{y}\right\|\leq\frac{2}{\mu}\nu\ .$ (22)
Putting (21) and (22) together, we have $f\left(\bar{y}\right)-f^{*}\leq{2}\nu^{2}/{\mu}$.
Summarizing two cases completes our proof. ∎∎
@@
1:Initialize: $x^{0}=\bar{y},\tilde{l}=f\left(\bar{y}\right)-\left(1+\beta\right)\Delta$, $\mathcal{P}=\left\{\bar{y}\right\}$
@@
8:  Compute $\bar{L}\left(t\right)$ (defined in (8)) based on $\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|$
@@
$\Psi_{t,f}=\left\{x:\psi_{\mathcal{P}_{t},f}\left(\hat{x}\right)\leq\tilde{l}\right\},\text{ where }\mathcal{P}_{t}=\left\{\underline{x}^{1,1},\ldots,\underline{x}^{1,m},\ldots,\underline{x}^{t-1,1},\ldots,\underline{x}^{t-1,m}\right\}\ ,$ (23)
@@
We prove this by induction. For base case $t=1$, we have $x^{2,0}=x^{1,m}$. Since $x^{1,0}=\bar{y}$, then $\bar{X}\left(t,0\right)$ in option I is eliminated. Hence, $x^{1,m}$ is the closest point to set $\left\{x:\psi_{\mathcal{P}_{1},f}\left(x\right)\leq\tilde{l}\right\}$, which implies that for any point satisfies $\hat{x}\in\Psi_{1,f}\cap X$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Suppose the claim (23) holds for $1,\ldots,t$, we aim to prove if every $\hat{x}\in\Psi_{t,f}\cap X$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Since $\hat{x}\in\Psi_{t-1,f}$ implies $\left\langle\hat{x}-x^{t,0},x^{t,0}-\bar{y}\right\rangle\geq 0$ by induction. Hence, the goal of our proof is equivalent to for any $\hat{x}\in X$, $\left\langle\hat{x}-x^{t,0},x^{t,0}-\bar{y}\right\rangle\geq 0$ and $\psi_{\mathcal{P}_{t}\backslash\mathcal{P}_{t-1},f}\left(\hat{x}\right)\leq\tilde{l}$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Note that the region
@@
Now, we are ready to prove the main result of Lemma 4.3. Since sequence $\left\{\left\|x^{t,i}-\bar{y}\right\|\right\}$ is non-decreasing by Proposition 3.1 and $\left\|x^{1,0}-\bar{y}\right\|=0$. Note that $x^{i,m}=x^{i+1,0}$ holds for any $i\in\left[t-1\right]$. Hence there exists a index pair $\left(\tilde{t},j\right)\in\left[t\right]\times\left[m\right]$ such that $\iota\in\left(\left\|x^{\tilde{t},j-1}-\bar{y}\right\|,\left\|x^{\tilde{t},j}-\bar{y}\right\|\right]$. We first denote $\mathcal{P}\left(\tilde{t},j\right)=\left\{\underline{x}^{\tilde{t},1},\ldots,\underline{x}^{\tilde{t},j}\right\}$, then the set $X\left(t,i\right)$ in Algorithm 1 can be rewritten as $X\left(\tilde{t},j\right)=\left\{x:\psi_{\mathcal{P}\left(\tilde{t},j\right),f}\left(x\right)\leq\tilde{l}\ ,\ \left\langle x-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle\geq 0\right\}$. Since $x^{\tilde{t},j}$ is the closest point to set $X\left(\tilde{t},j\right)$. Hence, $\forall\hat{x}\in X\left(\tilde{t},j\right)\cap X$, we have $\left\|\hat{x}-\bar{y}\right\|\geq\left\|x^{\tilde{t},j}-\bar{y}\right\|\geq\iota$. This implies that for any $\hat{x}\in\mathcal{B}\left(\bar{y};\iota\right)\cap X$, we have $\psi_{\mathcal{P}\left(\tilde{t},j\right),f}\left(\hat{x}\right)>\tilde{l}$ or $\left\langle\hat{x}-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle<0$. It follows from every $\hat{x}\in\Psi_{t,f}\cap X$ implies $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$ that $\left\langle\hat{x}-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle<0$ implies $\psi_{\mathcal{P}_{\tilde{t}},f}\left(\hat{x}\right)>\tilde{l}$. Hence, we complete the proof. ∎∎
@@
If $t\geq\sqrt{\left(2\iota_{\max}^{2}\bar{L}\left(t\right)+6\left(1+\beta\right)\Delta\right)/\beta\Delta}$, then we must have $\left\|x^{t,0}-\bar{y}\right\|\geq\iota_{\max}$, which implies this result. ∎∎
@@ 5 Restarted APEX for known $\mu^{*}$
    If $L_{\text{avg}}=\max_{s=1,\ldots}L\left(\bar{t}_{s}\right)$ (defined in (8)) of approximate Lipschitz constant $L_{t}$ (defined in (8)) across every outer loop, where $\bar{t}_{s}$ is the number of inner loop iterations consumed in each $s$-th outer loop, then to obtain an $\varepsilon$-optimal solution, $\left(f\left(x\right)-f^{*}\leq\varepsilon\right)$, the overall gradient evaluation calls is at most $O(1)\left(m\cdot\sqrt{{L_{\textrm{avg}}}/{\mu^{*}}}\log\left(\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}/{2\mu^{*}\varepsilon}\right)\right)$.
@@
$\bar{t}=\left\lceil\sqrt{\frac{6\theta+\frac{4L_{\text{avg}}\theta}{\mu}}{\left(2\theta-1\right)}}\right\rceil$ (24)
@@
$\begin{split}\omega_{\bar{t}}\left(\left(2\theta-1\right)\Delta_{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\theta\Delta_{s-1}&\overset{(a)}{<}\omega_{\bar{t}}\left(f\left(\hat{x}_{s}^{t}\right)-\tilde{l}^{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}_{s}^{1}\right)-\tilde{l}^{s-1}\right)\\ &\leq\bar{L}\left(\bar{t}\right)\left\|x^{N}-\bar{y}\right\|^{2}\ ,\end{split}$ (25)
@@
Hence, combining $\omega_{1}\left(1-\alpha_{1}/2\right)=3$, (25), $\bar{L}\left(\bar{t}\right)\leq L_{\text{avg}}$ and (18) gives
$\begin{split}\left\|x^{\bar{t}}-\bar{y}\right\|^{2}&\geq\frac{\omega_{\bar{t}}\left(\left(2\theta-1\right)\Delta_{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\theta\Delta_{s-1}}{L_{\text{avg}}}\geq\frac{\left(\omega_{\bar{t}}\left(2\theta-1\right)-6\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\\ &=\frac{\left(\frac{\left(\bar{t}+2\right)\left(\bar{t}+3\right)}{2}\left(2\theta-1\right)-3\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\geq\frac{\left(\frac{1}{2}\left(2\theta-1\right)\bar{t}^{2}-3\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\overset{(a)}{\geq}\frac{2\theta\Delta_{s-1}}{\mu^{*}}\end{split}\ ,$ (26)
@@
$\begin{split}\bar{L}\left(\bar{t}\right)&\leq\frac{3\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}_{s}^{t-1}\right)-\tilde{l}^{s-1}\right)\right)}{\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}_{s}^{t-1}\right)-\tilde{l}^{s-1}\right)\Big/\left(\alpha_{t}\cdot L_{t}\right)}\overset{(a)}{\leq}\frac{3\theta\Delta_{s-1}\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\alpha_{t}}{\left(2\theta-1\right)\Delta_{s-1}\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\Big/\left(\alpha_{t}\cdot L_{t}\right)}\ ,\\ &\overset{(b)}{\leq}\frac{24\theta\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{-1}}{\left(2\theta-1\right)\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{-1}L_{t}^{-1}}\end{split}$ (27)
where $(a)$ follows from $\theta\Delta_{s-1}\geq f\left(\hat{x}_{s}^{t}\right)-\tilde{l}^{s-1}\geq\left(2\theta-1\right)\Delta_{s-1}$ holds by $\omega_{t}\alpha_{t}\leq 8\alpha_{t}^{-1}$. If function $f$ is $\left(k,L\right)$-PWS function with $k\leq m$, then we have $L_{t}\leq L$. Hence, $\bar{L}\left(\bar{t}\right)=O(1)L$. Combining this result and complexity in Part 2 yields the result in Part 3. ∎∎

### v2

@@ 3 Accelerated Prox-level method for piecewise smooth optimization
Note that the optimal solution $x^{*}$ is always feasible for the subproblem if we set the level $\tilde{l}=f^{*}$. Hence, any feasible solution to the subproblem $x^{t}$ should satisfy $\left\|x^{t}-\bar{y}\right\|\leq\left\|x^{*}-\bar{y}\right\|$. ∎
One remaining question is that, even though $\bar{L}\left(t\right)$ (the average of $L_{t}$, defined in (3.2)) is computable in practice, it is still unclear how to bound the constant $\bar{L}\left(t\right)$ as it appears in Proposition 3.1 and Theorem 3.1. In the following Proposition 3.2, we show that value $\bar{L}\left(t\right)$ can be upper bounded by a weighted average of empirical Lipschitz smoothness constants $\tilde{L}_{t}$, where $\tilde{L}_{t}$ is the best Lipschitz constant between $\bar{x}^{t,r}$ and $\underline{x}^{t,l+1}$ with $r>l$ in $t$-th outer iteration.
@@
then $L_{t}\leq\tilde{L}_{t}$ ($L_{t}$ is defined in (3.2)). Furthermore, let the weighted sequence $\omega_{t}$ satisfy $\omega_{t}=\omega_{t-1}\left(1-{\alpha_{t}}/{2}\right)^{-1}$. Then $\bar{L}\left(N\right)$ (defined in (3.2)) admits the upper bound:
$\bar{L}\left(N\right)<\frac{3\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\right)}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)\Big/\left(\alpha_{t}\cdot\tilde{L}_{t}\right)}\ .$ (3.12)
@@
where $(a)$ holds by the definition of $L_{t}$ and $(b)$ follows from $f\left(\hat{x}^{t}\right)\leq f\left(\hat{x}^{t-1}\right)$ and the definition of $\mathcal{E}(t)$. Finally, since $L_{t}\leq\tilde{L}_{t}$, the stated upper bound follows. ∎
@@ 4 Normalized Wolfe certificate
Definition 3 (Normalized Wolfe certificate, W-certificate).
@@
$\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)=\frac{1}{\iota}\max_{x\in\mathcal{B}\left(\bar{y},\iota\right)\cap X}\left(\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right)\leq\nu\ ,\ \text{where}\ \psi_{\mathcal{P},f}\left(x\right)=\max_{z\in\mathcal{P}}\left\{\ell_{f}\left(x,z\right)\right\}\ .$ (4.1)
@@
$\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ (defined in (4.1)) is non-increasing with respect to $\iota$.
@@
$\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\left\|\tilde{x}-\bar{y}\right\|}\geq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\left\|x_{\iota 1}-\bar{y}\right\|}\ .$ (4.2)
Combining (4.2), $\left\|\bar{y}-\tilde{x}\right\|=\frac{\iota_{2}}{\iota_{1}}\left\|x_{\iota 1}-\bar{y}\right\|$ gives $\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\iota_{2}}\geq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\iota_{1}}$. By the definition of $x_{\iota 2}$, we have $\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 1}\right)}{\iota_{1}}\leq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(\tilde{x}\right)}{\iota_{2}}\leq\frac{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{\iota 2}\right)}{\iota_{2}}$. Hence, we complete our proof of that $\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ is non-increasing with respect to $\iota$. ∎
@@
$\begin{split}f\left(\bar{y}\right)-f^{*}&\leq\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x_{p}^{*}\right)\\ &\leq\max_{x\in\mathcal{B}\left(\bar{y},\left\|x_{p}^{*}-\bar{y}\right\|\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &=\left\|x_{p}^{*}-\bar{y}\right\|\mathcal{V}_{\mathcal{P},f}\left(\left\|x_{p}^{*}-\bar{y}\right\|;\bar{y}\right)\\ &\leq\left\|x_{p}^{*}-\bar{y}\right\|\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)\leq\left\|x_{p}^{*}-\bar{y}\right\|\nu\end{split}\ .$ (4.3)
@@
$\left\|x_{p}^{*}-\bar{y}\right\|\leq\frac{2}{\mu}\nu\ .$ (4.4)
Putting (4.3) and (4.4) together, we have $f\left(\bar{y}\right)-f^{*}\leq{2}\nu^{2}/{\mu}$.
Summarizing two cases completes our proof. ∎
@@
1:Initialize: $x^{0}=\bar{y},\hat{x}^{0}=\bar{y},\tilde{l}=f\left(\bar{y}\right)-\left(1+\beta\right)\Delta$, $\mathcal{P}=\left\{\bar{y}\right\}$
@@
8:  Compute $\bar{L}\left(t\right)$ (defined in (3.2)) based on $\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|$
@@
$\Psi_{t,f}=\left\{x:\psi_{\mathcal{P}_{t},f}\left(\hat{x}\right)\leq\tilde{l}\right\},\text{ where }\mathcal{P}_{t}=\left\{\underline{x}^{1,1},\ldots,\underline{x}^{1,m},\ldots,\underline{x}^{t-1,1},\ldots,\underline{x}^{t-1,m}\right\}\ ,$ (4.5)
@@
We prove this by induction. For base case $t=1$, we have $x^{2,0}=x^{1,m}$. Since $x^{1,0}=\bar{y}$, then $\bar{X}\left(t,0\right)$ in option I is eliminated. Hence, $x^{1,m}$ is the closest point to set $\left\{x:\psi_{\mathcal{P}_{1},f}\left(x\right)\leq\tilde{l}\right\}$, which implies that for any point satisfies $\hat{x}\in\Psi_{1,f}\cap X$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Suppose the claim (4.5) holds for $1,\ldots,t$, we aim to prove if every $\hat{x}\in\Psi_{t,f}\cap X$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Since $\hat{x}\in\Psi_{t-1,f}$ implies $\left\langle\hat{x}-x^{t,0},x^{t,0}-\bar{y}\right\rangle\geq 0$ by induction. Hence, the goal of our proof is equivalent to for any $\hat{x}\in X$, $\left\langle\hat{x}-x^{t,0},x^{t,0}-\bar{y}\right\rangle\geq 0$ and $\psi_{\mathcal{P}_{t}\backslash\mathcal{P}_{t-1},f}\left(\hat{x}\right)\leq\tilde{l}$, we have $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$. Note that the region
@@
Now, we are ready to prove the main result of Lemma 4.3. Since sequence $\left\{\left\|x^{t,i}-\bar{y}\right\|\right\}$ is non-decreasing by Proposition 3.1 and $\left\|x^{1,0}-\bar{y}\right\|=0$. Note that $x^{i,m}=x^{i+1,0}$ holds for any $i\in\left[t-1\right]$. Hence there exists a index pair $\left(\tilde{t},j\right)\in\left[t\right]\times\left[m\right]$ such that $\iota\in\left(\left\|x^{\tilde{t},j-1}-\bar{y}\right\|,\left\|x^{\tilde{t},j}-\bar{y}\right\|\right]$. We first denote $\mathcal{P}\left(\tilde{t},j\right)=\left\{\underline{x}^{\tilde{t},1},\ldots,\underline{x}^{\tilde{t},j}\right\}$, then the set $X\left(t,i\right)$ in Algorithm 1 can be rewritten as $X\left(\tilde{t},j\right)=\left\{x:\psi_{\mathcal{P}\left(\tilde{t},j\right),f}\left(x\right)\leq\tilde{l}\ ,\ \left\langle x-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle\geq 0\right\}$. Since $x^{\tilde{t},j}$ is the closest point to set $X\left(\tilde{t},j\right)$. Hence, $\forall\hat{x}\in X\left(\tilde{t},j\right)\cap X$, we have $\left\|\hat{x}-\bar{y}\right\|\geq\left\|x^{\tilde{t},j}-\bar{y}\right\|\geq\iota$. This implies that for any $\hat{x}\in\mathcal{B}\left(\bar{y};\iota\right)\cap X$, we have $\psi_{\mathcal{P}\left(\tilde{t},j\right),f}\left(\hat{x}\right)>\tilde{l}$ or $\left\langle\hat{x}-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle<0$. It follows from every $\hat{x}\in\Psi_{t,f}\cap X$ implies $\left\langle\hat{x}-x^{t+1,0},x^{t+1,0}-\bar{y}\right\rangle\geq 0$ that $\left\langle\hat{x}-x^{\tilde{t},0},x^{\tilde{t},0}-\bar{y}\right\rangle<0$ implies $\psi_{\mathcal{P}_{\tilde{t}},f}\left(\hat{x}\right)>\tilde{l}$. Hence, we complete the proof. ∎
@@
If $t\geq\sqrt{\left(2\iota_{\max}^{2}\bar{L}\left(t\right)+6\left(1+\beta\right)\Delta\right)/\beta\Delta}$, then we must have $\left\|x^{t,0}-\bar{y}\right\|\geq\iota_{\max}$, which implies this result. ∎
@@ 5 Restarted APEX for known $\mu^{*}$
    If $L_{\text{avg}}=\max_{s=1,\ldots}L\left(\bar{t}_{s}\right)$ (defined in (3.2)) of approximate Lipschitz constant $L_{t}$ (defined in (3.2)) across every outer loop, where $\bar{t}_{s}$ is the number of inner loop iterations consumed in each $s$-th outer loop, then to obtain an $\varepsilon$-optimal solution, $\left(f\left(x\right)-f^{*}\leq\varepsilon\right)$, the overall gradient evaluation calls is at most $O(1)\left(m\cdot\sqrt{{L_{\textrm{avg}}}/{\mu^{*}}}\log\left(\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}/{2\mu^{*}\varepsilon}\right)\right)$.
@@
$\bar{t}=\left\lceil\sqrt{\frac{6\theta+{4L_{\text{avg}}\theta}/{\mu}}{\left(2\theta-1\right)}}\right\rceil$ (5.1)
@@
$\begin{split}\omega_{\bar{t}}\left(\left(2\theta-1\right)\Delta_{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\theta\Delta_{s-1}&\overset{(a)}{<}\omega_{\bar{t}}\left(f\left(\hat{x}_{s}^{t}\right)-\tilde{l}^{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}_{s}^{1}\right)-\tilde{l}^{s-1}\right)\\ &\leq\bar{L}\left(\bar{t}\right)\left\|x^{N}-\bar{y}\right\|^{2}\ ,\end{split}$ (5.2)
@@
Hence, combining $\omega_{1}\left(1-\alpha_{1}/2\right)=3$, (5.2), $\bar{L}\left(\bar{t}\right)\leq L_{\text{avg}}$ and (3.12) gives
$\begin{split}\left\|x^{\bar{t}}-\bar{y}\right\|^{2}&\geq\frac{\omega_{\bar{t}}\left(\left(2\theta-1\right)\Delta_{s-1}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\theta\Delta_{s-1}}{L_{\text{avg}}}\geq\frac{\left(\omega_{\bar{t}}\left(2\theta-1\right)-6\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\\ &=\frac{\left(\frac{\left(\bar{t}+2\right)\left(\bar{t}+3\right)}{2}\left(2\theta-1\right)-3\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\geq\frac{\left(\frac{1}{2}\left(2\theta-1\right)\bar{t}^{2}-3\theta\right)\Delta_{s-1}}{L_{\text{avg}}}\overset{(a)}{\geq}\frac{2\theta\Delta_{s-1}}{\mu^{*}}\end{split}\ ,$ (5.3)
@@
$\begin{split}\bar{L}\left(\bar{t}\right)&\leq\frac{3\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\left(\alpha_{t}\left(f\left(\hat{x}_{s}^{t-1}\right)-\tilde{l}^{s-1}\right)\right)}{\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\left(f\left(\hat{x}_{s}^{t-1}\right)-\tilde{l}^{s-1}\right)\Big/\left(\alpha_{t}\cdot L_{t}\right)}\overset{(a)}{\leq}\frac{3\theta\Delta_{s-1}\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\omega_{t}\alpha_{t}}{\left(2\theta-1\right)\Delta_{s-1}\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\Big/\left(\alpha_{t}\cdot L_{t}\right)}\ ,\\ &\overset{(b)}{\leq}\frac{24\theta\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{-1}}{\left(2\theta-1\right)\sum_{t=1}^{\bar{t}}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{-1}L_{t}^{-1}}\end{split}$ (5.4)
where $(a)$ follows from $\theta\Delta_{s-1}\geq f\left(\hat{x}_{s}^{t}\right)-\tilde{l}^{s-1}\geq\left(2\theta-1\right)\Delta_{s-1}$ holds by $\omega_{t}\alpha_{t}\leq 8\alpha_{t}^{-1}$. If function $f$ is $\left(k,L\right)$-PWS function with $k\leq m$, then we have $L_{t}\leq L$. Hence, $\bar{L}\left(\bar{t}\right)=O(1)L$. Combining this result and complexity in Part 2 yields the result in Part 3. ∎